THE ROTATIONAL SPECTRUM OF HCl^{+}

\underline{HARSHAL GUPTA}, B. J. DROUIN, J. C. PEARSON, Jet Propulsion Laboratory, California Institute of Technology^{a}, Pasadena, CA 91109.

The rotational spectrum of the radical ion HCl^{+} has been detected at high resolution in the laboratory, supporting conclusively the identification of this ion with the Herschel Space Observatory’s Heterodyne Instrument for the Far-Infrared (HIFI) in diffuse clouds toward the Galactic star-forming regions W31C and W49N. Three rotational transitions, one in the ground state ^{2}Π_{3/2} ladder and two in the ^{2}Π_{1/2} ladder (643 cm^{−1} above ground), were observed in a microwave discharge of He and HCl. Well-resolved chlorine hyperfine structure and Λ-doubling, and the detection of lines of H^{37}Cl^{+} at precisely the expected isotopic shift, provide conclusive evidence for the laboratory identification. The detection of rotational transitions in the ^{2}Π_{1/2} ladder of HCl^{+} for the first time allows an experimental determination of the individual hyperfine coupling constants of chlorine, and yields a precise value of eQq_{2}. The spectroscopic constants determined by fitting a Hamiltonian simultaneously to our data and more than 8000 optical transitions are so precise, that they allow calculation of the frequencies of ^{2}Π_{3/2} J = 5/2 − 3/2 transition observed by HIFI to within 0.2 km s^{−1}, and indeed, those of the strongest rotational transitions of HCl^{+} below 7.5 THz to better than 1 km s^{−1}.

aA part of this work was performed at the Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration. Copyright 2012\begin{matrix}c \\ ◯ \end{matrix} California Institute of Technology. All rights reserved.